What is the x - coordinate of the vertex of y = 4 x^2 + 24x - 7
Answer by lwsshak3(11628)

What is the x - coordinate of the vertex of y = 4 x^2 + 24x - 7
vertex form of parabola: y=A(x-h)^2+k, vertex:(h,k)
y = 4 x^2 + 24x - 7
complete the square:
y=4 (x^2 +6x+9)-36 - 7
y=4 (x+3)^2-43
This is an equation of a parabola that opens up with vertex at (-3,-43)
x-coordinate of vertex=-3